How many liters of water must be evaporated from 10 L of a 40% salt solution to produce a 50% solution?
Found 4 solutions by ankor@dixie-net.com, amalm06, ikleyn, greenestamps:
Answer by ankor@dixie-net.com(22740)

How many liters of water must be evaporated from 10 L of a 40% salt solution to produce a 50% solution?
:
Let x = amt of water to evaporated
:
Write an amt of salt equation (amt of salt remains unchanged)
.4(10) = .5(10-x)
4 = 5 - .5x
.5x = 5 - 4
.5x = 1
x = 2 Liters to be evaporated

Answer by amalm06(224)

The problem can be solved using the method of alligation.

Let x denote water and y denote the 40% solution

Then 50-40=0 and 50-0=50

The ratio of x/y is 10/50 or 1/5

Therefore, the amount of water that must be evaporated is (10/50)(10)=2 L (Answer)

Answer by ikleyn(52786)

.

For your info:

    No salt solution with the concentration above 27% - 28%.


    It does not exist and can not exist.


    See this Wikipedia article    https://en.wikipedia.org/wiki/Saline_water


Answer by greenestamps(13200)



Ignoring the scientific fact that these percentages of salt solution are not possible, and treating this purely as a mathematical exercise, note that the amount of salt in 10 L of 40% salt is 4L.

You want that same amount of salt to be 50% of the final mixture; that means the final mixture must be 8 L.

Since you started with 10L and ended with 8L, the amount of water evaporated was 2L.